The average of four consecutive eve integers is 17. Find the integers.
Found 2 solutions by user_dude2008, mananth:
Answer by user_dude2008(1862)

(x+x+2+x+4+x+6)/4=17


(4x+12)/4=17


4x+12=68


4x=68-12

4x=56

x=56/4

x=14

numbers are: 14, 16, 18, 20

Answer by mananth(16949)

let the even integers be x,x+2,x+4,x+6
sum =4x+10
..
average = 17
17*4= 68 = sum
4x+12 = 68
add-12
4x+12-12=68-12
4x=56
/4
x=14
...
14,16,18,20
...
m.ananth@hotmail.ca